# Adversarial thermodynamics

### v1

@@ Appendix B General protocol for adversarial work extraction
In this section, we derive the work formula for a general adversarial engine. Let $x$ be a microscopic state that is not necessarily binary. Suppose that the initial distribution is given by $P_{X}(x)$ and the initial energy level is given by $E_{X}^{B}(x)$. We relate the energy level to Bob’s distribution by $Q_{X}^{B}(x)=e^{-E_{X}^{B}(x)}/Z_{X}^{B}$ with $Z_{X}^{B}=\sum_{x}e^{-E_{X}^{B}(x)}$ being the normalization factor (partition function). Note that we set $k_{B}T$ to unity throughout this section.
@@
Let us first remember the optimal work extraction protocol at the level of ensemble average: (i) Alice quenches the energy level to $E_{X}(x)$ that is defined through $P_{X}(x)=e^{-E_{X}(x)}/Z_{X}$ with $Z_{X}=\sum_{x}e^{-E_{X}(x)}$; (ii) Alice moves the energy level from $E_{X}(x)$ to $E_{X}^{B}(x)$ quasi-statically and isothermally. It is well known that Alice extracts the work $D(P_{X}\|Q_{X}^{B})k_{B}T$ on average by this protocol [1, 3].
@@
Now we turn to the adversarial scenario. The protocol that Alice performs is: (i) Alice quenches the energy level to $E_{X}^{A}(x)$ that is defined through $Q_{X}^{A}(x)=e^{-E_{X}^{A}(x)}/Z_{X}^{A}$ with $Z_{X}^{A}=\sum_{x}e^{-E_{X}^{A}(x)}$ for her choice of $Q_{X}^{A}$; (i’) Alice lets the system thermalized; (ii) Alice moves the energy level from $E_{X}^{A}(x)$ to $E_{X}^{B}(x)$ quasi-statically and isothermally. The extracted work in step (i) is given by $E_{X}^{B}(x)-E_{X}^{A}(x)$, while in step (ii) $\ln Z_{X}^{B}-\ln Z_{X}^{A}$ (the change in the equilibrium free energy). We thus obtain
$w_{X}(x)=E_{X}^{B}(x)-E_{X}^{A}(x)+\ln Z_{X}^{B}-\ln Z_{X}^{A}=\ln\left(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\right),$ (24)
@@
which reproduces Eq. (12). By taking the average with respect to $P_{X}$, Eq. (1) is also reproduced.
Appendix C Work extraction and the method of types
This appendix summarizes the application of the method of types to the adversarial Szilard engine, following the general approach detailed in concurrent work [17]. The method of types [32] provides a powerful framework for analyzing finite-size fluctuations in information-theoretic tasks.
For a sequence of $n$ rounds, the empirical distribution (type) $\lambda_{x^{n}}$ occurs with probability approximately $\exp(-nD(\lambda_{x^{n}}\|P_{X}))$. The work extracted when the type is $\lambda_{x^{n}}$ is $nD(\lambda_{x^{n}}\|Q_{X}^{B})k_{B}T$ when Alice sets $Q_{X}^{A}=\lambda_{x^{n}}$.
The optimization problem becomes: maximize the reward $D(\lambda_{x^{n}}\|Q_{X}^{B})$ subject to the constraint that the probability of success exceeds $\epsilon$, i.e., $D(\lambda_{x^{n}}\|P_{X})\leq\frac{1}{n}\ln(1/\epsilon)$. Using Lagrange multipliers, one obtains the family of optimal strategies:
$Q_{X}^{A*}(x)=\frac{P_{X}(x)^{\mu}Q_{X}^{B}(x)^{1-\mu}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\mu}Q_{X}^{B}(x^{\prime})^{1-\mu}},$ (25)
where $\mu\in[0,1]$ is determined by the constraint. Substituting back gives the work bound:
$W_{n}\geq nD_{\mu}(P_{X}\|Q_{X}^{B})k_{B}T+\frac{\mu}{1-\mu}\ln\epsilon.$ (26)
For a comprehensive treatment of the method of types in adversarial scenarios, including the connection to hypothesis testing and utility theory, we refer to [17, 21].
Appendix D Maximisation of the utility function
In this section, we compute the maximiser of the utility function for risk parameter $r$. Recall now that the extracted work from the adversarial Szilard engine for a single round can be described by the random variable
$w_{X}(x)=\ln\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)k_{B}T$ (27)
It follows using eq. 2 that the utility of outcome $x$ is
$u_{r}(w_{x})=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (28)
By the expected utility hypothesis, a rational gambler will set their bet $Q_{X}^{A}$ as to maximise
$\sum_{x}P_{X}(x)u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (29)
@@ Appendix D Maximisation of the utility function
Note that the first term is a constant, so we need only focus on the second term. For $r>0$ (risk-aversion), the second term is negative, whilst if $r<0$ (risk-seeking), the second term is positive. This means that for a risk-averse individual, we minimise
$\min_{Q_{X}^{A}}\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}$ (30)
which is equivalent to minimising
$\min_{Q_{X}^{A}}\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})$ (31)
whilst for a risk-seeking individual, we maximise
$\max_{Q_{X}^{A}}\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}$ (32)
which is equivalent to maximising
$\max_{Q_{X}^{A}}\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})$ (33)
@@
In order to find the minimiser (resp. maximiser), we will use a lemma proved in [26], which states that
$\ln(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r})=-r\big(D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})-D_{1+r}(G^{r}_{X}||Q_{X}^{A})\big)$ (34)
@@
$G_{X}^{r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (35)
The maximiser is therefore
$Q_{X}^{A,r}=\frac{P_{X}(x)^{\frac{1}{1+r}}Q_{X}^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q_{X}^{B}(x^{\prime})^{\frac{r}{1+r}}}$ (36)
Appendix E Calculation of the certainty equivalent
At the beginning of this section, we said that two equivalent ways of characterising risk preferences were through the concavity/convexity of the utility function and the certainty equivalent. As shown in [16], the certainty equivalent has a clear information-theoretic interpretation in a thermodynamic context. We will also find it useful here to calculate the certainty equivalent of work for each value of the risk aversion parameter $r$ below.
Recall that, for an agent exhibiting risk aversion, the certainty equivalent is the guaranteed amount of money that the individual would accept instead of taking a risky bet with different payoff. We then say that an individual with utility function $u(.)$ is risk averse if and only if their certainty equivalent is lower than the expected value of the risky lottery. The mathematical definition of the certainty equivalent is the number $W_{CE}\in\mathbb{R}$ such that $u_{r}(W_{CE})=E(u_{r}(w_{X}(x)))$, where $E$ denotes expected value. Since the utility of outcome $x$ is
$u_{r}(w_{X}(x))=\frac{1}{r}\bigg(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (37)
we must calculate the number $W_{CE}$ $u_{r}(W_{CE})=E(u_{r}(x))$, where $E$ denotes expected value, we therefore calculate
$W_{CE}=u_{r}^{-1}\big(E(u_{r}(w))\big)\\ =u_{r}^{-1}\bigg(E\bigg(\frac{1}{r}(1-\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)\bigg)\\ =u_{r}^{-1}\bigg(\frac{1}{r}\bigg(1-\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)\bigg)\\ =-\frac{1}{r}\ln\bigg(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (38)
where in the last line we have used that the inverse function of $u_{r}(x)$ is $u_{r}^{-1}(x)=-\frac{1}{r}\ln(1-rx)$. Since
$W_{CE}=-\frac{1}{r}\ln\bigg(\sum_{x}P_{X}(x)\bigg(\frac{Q_{X}^{A}(x)}{Q_{X}^{B}(x)}\bigg)^{-r}\bigg)$ (39)
which equals
$D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})-D_{1+r}(G^{r}_{X}||Q_{X}^{A})$ (40)
by the result of [26], we finally conclude that
$W_{CE}=D_{\frac{1}{1+r}}(P_{X}||Q^{B}_{X})k_{B}T$ (41)
when the expression for $Q_{X}^{A,r}$ is substituted into the expression for the certainty equivalent.
@@ Appendix E Calculation of the certainty equivalent
In addition, the work certainty equivalent is smaller than $D_{1}$ (the non-equilibrium free energy) for a risk averse player, and equal to $D_{1}$ for a risk neutral player. In particular, for an extremely risk averse player ($r\to\infty$), the certainty equivalent tends to $D_{0}(P_{X}||Q_{X}^{B})$, which becomes zero when both distributions have the same support.
Appendix F Extracted work as a function of strategy
In section II, we showed that for a given strategy of Alice $Q_{X}^{A}$, the average work extracted in the adversarial Szilard engine is given by
$\mathbb{E}[W]=\big(D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A}_{X})\big)k_{B}T$ (42)
Our goal is to express this for the optimal strategy $Q_{X}^{A}$ which from now on we write as $Q^{A,r}_{X}$, to emphasise dependence on the risk parameter of the CARA utility function. We substitute the explicit form of the optimal strategy into the second term.
The optimal strategy for a given risk parameter $r$ is:
$Q^{A,r}X(x)=\frac{P_{X}(x)^{\frac{1}{1+r}}Q^{B}_{X}(x)^{\frac{r}{1+r}}}{Z},$ (43)
@@ Appendix F Extracted work as a function of strategy
$Z(\alpha)=\sum_{x^{\prime}}P_{X}(x^{\prime})^{\frac{1}{1+r}}Q^{B}_{X}(x^{\prime})^{\frac{r}{1+r}}$ (44)
Let $\alpha=\frac{1}{1+r}$ for notational clarity. Since the first divergence is independent of the strategy $Q^{A,r}_{X}$, we compute the second divergence term, $D(P_{X}||Q^{A,r}_{X})$:
$D(P_{X}||Q^{A,r}_{X})$ $=\sum_{x}P_{X}(x)\ln\frac{P_{X}(x)}{Q^{A,r}_{X}(x)}$ (45)
$=\sum_{x}P_{X}(x)\left[\ln P_{X}(x)-\ln\left(\frac{P_{X}(x)^{\alpha}Q^{B}_{X}(x)^{1-\alpha}}{Z}\right)\right]$ (46)
$=\sum_{x}P_{X}(x)\left[\ln P_{X}(x)-\alpha\ln P_{X}(x)-(1-\alpha)\ln Q^{B}_{X}(x)+\ln Z\right]$ (47)
$=(1-\alpha)\sum_{x}P_{X}(x)\ln\frac{P_{X}(x)}{Q^{B}_{X}(x)}+\ln Z$ (48)
$=(1-\alpha)D(P_{X}||Q^{B}_{X})+\ln Z(\alpha)$ (49)
The normalization constant $Z$ is related to the sum that defines the Rényi divergence. Recall its definition:
$D_{\alpha}(P_{X}||Q^{B}_{X})=\frac{1}{\alpha-1}\ln\sum_{x}P_{X}(x)^{\alpha}Q^{B}_{X}(x)^{1-\alpha}=\frac{1}{\alpha-1}\ln Z$ (50)
So that $\ln Z$:
$\ln Z(\alpha)=(\alpha-1)D_{\alpha}(P_{X}||Q^{B}_{X})$ (51)
@@
$D(P_{X}||Q^{A,r}X)$ $=(1-\alpha)D(P_{X}||Q^{B}_{X})+(\alpha-1)D\alpha(P_{X}||Q^{B}_{X})$ (52)
$=(1-\alpha)\left[D(P_{X}||Q^{B}_{X})-D_{\alpha}(P_{X}||Q^{B}_{X})\right]$ (53)
@@
$\mathbb{E})[W]=D(P_{X}||Q^{B}_{X})-D(P_{X}||Q^{A,r}X)\\ =D(P_{X}||Q^{B}_{X})-(1-\alpha)\left[D(P_{X}||Q^{B}_{X})-D_{\alpha}(P_{X}||Q^{B}_{X})\right]\\ =D(P_{X}||Q^{B}_{X})-(1-\alpha)D(P_{X}||Q^{B}_{X})+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})\\ =\alpha D(P_{X}||Q^{B}_{X})+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})$ (54)
@@
Thus, the expected work extracted for the optimal strategy corresponding to risk parameter $r$ is:
$\boxed{\mathbb{E}[W]=\alpha D(P_{X}||Q^{B}_{X})k_{B}T+(1-\alpha)D_{\alpha}(P_{X}||Q^{B}_{X})k_{B}T,\quad\text{where}\quad\alpha=\frac{1}{1+r}}$ (55)
Appendix G Conditions for rationality for r<1r<-1
@@
In section eq. 41 we showed that the certainty equivalent for work extraction for a rational agent with CARA utility function is a Rényi divergence:
$W_{CE}=D_{\alpha}(P_{X}||Q_{X}^{B})k_{B}T\quad\text{where}\quad\alpha=\frac{1}{1+r}$ (56)
The work extracted in a specific outcome $x$ is:
$W(x)=\ln\left(\frac{Q_{X}^{A,r}(x)}{Q_{X}^{B}(x)}\right)k_{B}T$
@@
where $Q_{X}^{A,r}$ is the optimal strategy for a given value of the risk parameter $r$.
The violation of rationality (First-Order Stochastic Dominance) occurs if the certainty equivalent is more negative than the worst-case work output:
$W_{CE}<\min_{x}W(x)$
Substituting in the expression for the certainty equivalent, we get:
$D_{\alpha}(P_{X}||Q_{X}^{B})<\min_{x}\ln\left(\frac{Q_{X}^{A,r}(x)}{Q_{X}^{B}(x)}\right)$
@@
This is the general condition. We can now use the specific form of $Q_{X}^{A,r}$ to simplify the right-hand side. Recall:
$Q_{X}^{A,r}(x)=\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{1-\alpha}}{\sum_{x^{\prime}}P_{X}(x^{\prime})^{\alpha}Q_{X}^{B}(x^{\prime})^{1-\alpha}}$
Defining $Z(\alpha)=\sum_{x^{\prime}}P_{X}(x^{\prime})^{\alpha}Q_{X}^{B}(x^{\prime})^{1-\alpha}$ as the normalisation constant and substituting into the work expression:
$W(x)=\ln\left(\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{1-\alpha}}{Z(\alpha)Q_{X}^{B}(x)}\right)=\ln\left(\frac{P_{X}(x)^{\alpha}Q_{X}^{B}(x)^{-\alpha}}{Z(\alpha)}\right)=\alpha\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
@@
To find the minimum value of this expression, note that the term $\ln Z(\alpha)$ is independent of $x$. Therefore:
$\min_{x}W(x)=\alpha\cdot\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
The normalisation constant $\ln Z(\alpha)$ is directly related to the Rényi divergence by:
$D_{\alpha}(P_{X}||Q_{X}^{B})=\frac{1}{\alpha-1}\ln Z(\alpha)$
Now, we can substitute back into our violation condition (Eq. 1):
$D_{\alpha}(P_{X}||Q_{X}^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)-\ln Z(\alpha)$
Substitute $D_{\alpha}$ and rearrange:
$\frac{1}{\alpha-1}\ln Z(\alpha)<\alpha\cdot M-\ln Z(\alpha)\quad\text{where}\quad M=\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)$
$\ln Z(\alpha)\left(\frac{1}{\alpha-1}+1\right)<\alpha M$
$\ln Z(\alpha)\left(\frac{1+(\alpha-1)}{\alpha-1}\right)<\alpha M$
$\frac{\ln Z(\alpha)\cdot\alpha}{\alpha-1}<\alpha M$
For $\alpha<0$ (which is true for r<1r<-1, the case we are interested in), dividing both sides by $\alpha$ flips the sign of the inequality:
$\frac{\ln Z(\alpha)}{\alpha-1}>M$
Since the left-hand side is exactly the Rényi divergence $D_{\alpha}(P_{X}||Q_{X}^{B})$, the condition for a violation of rationality simplifies to:
$\boxed{D_{\alpha}(P_{X}||Q_{X}^{B})>\max_{x}\ln\left(\frac{P_{X}(x)}{Q_{X}^{B}(x)}\right)}$
since this never holds, rationality is never violated.

### v2

@@ Appendix B General protocol for adversarial work extraction
In this appendix, we derive the work formula for a general adversarial engine. Let $x$ be a microscopic state that is not necessarily binary. Suppose that the initial distribution is given by $P(x)$ and the initial energy level is given by $E^{B}(x)$. We relate the energy level to Bob’s distribution by $Q^{B}(x)=e^{-E^{B}(x)}/Z^{B}$ with $Z^{B}=\sum_{x}e^{-E^{B}(x)}$ being the normalization factor (partition function). Note that we set $k_{B}T$ to unity throughout this section.
@@
Let us first remember the optimal work extraction protocol at the level of ensemble average: (i) Alice quenches the energy level to $E(x)$ that is defined through $P(x)=e^{-E(x)}/Z$ with $Z=\sum_{x}e^{-E(x)}$; (ii) Alice moves the energy level from $E(x)$ to $E^{B}(x)$ quasi-statically and isothermally. It is well known that Alice extracts the work $D(P\|Q^{B})$ on average by this protocol Sagawa (2022); Parrondo et al. (2015).
@@
Now we turn to the adversarial scenario. The protocol that Alice performs is: (i) Alice quenches the energy level to $E^{A}(x)$ that is defined through $Q^{A}(x)=e^{-E^{A}(x)}/Z^{A}$ with $Z^{A}=\sum_{x}e^{-E^{A}(x)}$ for her choice of $Q^{A}$; (i’) Alice lets the system thermalize; (ii) Alice moves the energy level from $E^{A}(x)$ to $E^{B}(x)$ quasi-statically and isothermally. The extracted work in step (i) is given by $E^{B}(x)-E^{A}(x)$, while in step (ii) $\ln Z^{B}-\ln Z^{A}$ (the change in the equilibrium free energy). We thus obtain
$w(x)=E^{B}(x)-E^{A}(x)+\ln Z^{B}-\ln Z^{A}=\ln\left(\frac{Q^{A}(x)}{Q^{B}(x)}\right),$ (29)
@@
which reproduces Eq. (5). By taking the average with respect to $P$, Eq. (3) is also reproduced.
Appendix C maximization of the utility function
@@
In this appendix, we compute the maximizer of the utility function for risk parameter $r$. Recall that a rational gambler will set their bet $Q^{A}$ as to maximize Eq. (7). Note that the first term on the right-hand side is a constant, so we need only focus on the second term. For $r>0$ (risk-aversion), the second term is negative, whilst if $r<0$ (risk-seeking), the second term is positive. This means that for a risk-averse individual, we minimize
$\min_{Q^{A}}\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r},$ (30)
which is equivalent to minimizing
$\min_{Q^{A}}\ln(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}).$ (31)
On the other hand, for a risk-seeking individual, we maximize
$\max_{Q^{A}}\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r},$ (32)
which is equivalent to maximizing
$\max_{Q^{A}}\ln(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}).$ (33)
@@
In order to find the minimizer (resp. maximizer), we will use a lemma proved in Bleuler et al. (2020), which states that
$\ln(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r})=-r\big(D_{\frac{1}{1+r}}(P||Q^{B})-D_{1+r}(G^{r}||Q^{A})\big),$ (34)
@@
$G^{r}=\frac{P(x)^{\frac{1}{1+r}}Q^{B}(x)^{\frac{r}{1+r}}}{\sum_{x^{\prime}}P(x^{\prime})^{\frac{1}{1+r}}Q^{B}(x^{\prime})^{\frac{r}{1+r}}}.$ (35)
The minimizer (resp. maximizer) is therefore given by Eq. (8).
Appendix D Calculation of the certainty equivalent
Recall that two equivalent ways of characterizing risk preferences were through the concavity/convexity of the utility function and the certainty equivalent. As shown in Ducuara et al. (2023), the certainty equivalent has a clear information-theoretic interpretation in a thermodynamic context. We will also find it useful here to calculate the certainty equivalent of work for each value of the risk aversion parameter $r$ below.
Since the utility of outcome $x$ is given by Eq. (6), we calculate
$W_{CE}$ $=u_{r}^{-1}\big(\mathbb{E}(u_{r}(w(x)))\big)$ (36)
$=u_{r}^{-1}\bigg(\mathbb{E}\bigg(\frac{1}{r}(1-\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg)\bigg)$ (37)
$=u_{r}^{-1}\bigg(\frac{1}{r}\bigg(1-\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg)\bigg)$ (38)
$=-\frac{1}{r}\ln\bigg(\sum_{x}P(x)\bigg(\frac{Q^{A}(x)}{Q^{B}(x)}\bigg)^{-r}\bigg),$ (39)
where in the last line we have used that the inverse function of $u_{r}(x)$ is $u_{r}^{-1}(x)=-\frac{1}{r}\ln(1-rx)$. Again from the result of Bleuler et al. (2020), we obtain
$W_{CE}=D_{\frac{1}{1+r}}(P||Q^{B})-D_{1+r}(G^{r}||Q^{A}).$ (40)
We finally conclude Eq. (11), when the expression for $Q^{A,r}$ is substituted into the expression for the certainty equivalent.
@@ Appendix D Calculation of the certainty equivalent
In addition, the work certainty equivalent is smaller than $D_{1}$ (the nonequilibrium free energy) for a risk averse player, and equal to $D_{1}$ for a risk neutral player. In particular, for an extremely risk averse player ($r\to\infty$), the certainty equivalent tends to $D_{0}(P||Q^{B})$, which becomes zero when both distributions have the same support.
Appendix E Extracted work as a function of strategy
Our goal of this Appendix is to express Eq. (3) for the optimal strategy $Q^{A}$ which from now on we write as $Q^{A,r}$, to emphasize dependence on the risk parameter of the CARA utility function. We substitute the explicit form of the optimal strategy into the second term.
Let $\alpha=\frac{1}{1+r}$ for notational clarity. The optimal strategy for a given risk parameter $r$ is given by
$Q^{A,r}(x)=\frac{P(x)^{\alpha}Q^{B}(x)^{1-\alpha}}{Z(\alpha)},$ (41)
@@ Appendix E Extracted work as a function of strategy
$Z(\alpha)=\sum_{x^{\prime}}P(x^{\prime})^{\alpha}Q^{B}(x^{\prime})^{1-\alpha}.$ (42)
Since the first divergence is independent of the strategy $Q^{A,r}$, we compute the second divergence term:
$D(P||Q^{A,r})$ $=\sum_{x}P(x)\ln\frac{P(x)}{Q^{A,r}(x)}$ (43)
$=\sum_{x}P(x)\left[\ln P(x)-\ln\left(\frac{P(x)^{\alpha}Q^{B}(x)^{1-\alpha}}{Z(\alpha)}\right)\right]$ (44)
$=\sum_{x}P(x)\left[\ln P(x)-\alpha\ln P(x)-(1-\alpha)\ln Q^{B}(x)+\ln Z(\alpha)\right]$ (45)
$=(1-\alpha)\sum_{x}P(x)\ln\frac{P(x)}{Q^{B}(x)}+\ln Z(\alpha)$ (46)
$=(1-\alpha)D(P||Q^{B})+\ln Z(\alpha).$ (47)
From Eq. (10), the normalization constant $Z(\alpha)$ is related to the Rényi divergence:
$\ln Z(\alpha)=(\alpha-1)D_{\alpha}(P||Q^{B}).$ (48)
@@
$D(P||Q^{A,r})$ $=(1-\alpha)D(P||Q^{B})+(\alpha-1)D_{\alpha}(P||Q^{B})$ (49)
$=(1-\alpha)\left[D(P||Q^{B})-D_{\alpha}(P||Q^{B})\right].$ (50)
@@
$D(P||Q^{B})-D(P||Q^{A,r})$ $=D(P||Q^{B})-(1-\alpha)\left[D(P||Q^{B})-D_{\alpha}(P||Q^{B})\right]$ (51)
$=D(P||Q^{B})-(1-\alpha)D(P||Q^{B})+(1-\alpha)D_{\alpha}(P||Q^{B})$ (52)
$=\alpha D(P||Q^{B})+(1-\alpha)D_{\alpha}(P||Q^{B}).$ (53)
@@
Thus, the expected work extracted for the optimal strategy corresponding to risk parameter $r$ is given by Eq. (9).
Appendix F Conditions for rationality for r<1r<-1
@@
In Eq. (11), we showed that the certainty equivalent for work extraction for a rational agent with CARA utility function is a Rényi divergence. The work extracted in a specific outcome $x$ is
$w(x)=k_{B}T\ln\left(\frac{Q^{A,r}(x)}{Q^{B}(x)}\right),$ (54)
@@
where $Q^{A,r}$ is the optimal strategy for a given value of the risk parameter $r$. Substituting this to the certainty equivalent, we get
$D_{\alpha}(P||Q^{B})<\min_{x}\ln\left(\frac{Q^{A,r}(x)}{Q^{B}(x)}\right).$ (55)
@@
This is the general condition. We can now use the specific form of $Q^{A,r}$ to simplify the right-hand side. From Eq. (42) and Eq. (48), we obtain
$w(x)=\ln\left(\frac{P(x)^{\alpha}Q^{B}(x)^{1-\alpha}}{Z(\alpha)Q^{B}(x)}\right)=\ln\left(\frac{P(x)^{\alpha}Q^{B}(x)^{-\alpha}}{Z(\alpha)}\right)=\alpha\ln\left(\frac{P(x)}{Q^{B}(x)}\right)-(\alpha-1)D_{\alpha}(P||Q^{B}).$ (56)
@@
To find the minimum value of this expression, recall that $\alpha<0$ for r<1r<-1. Therefore,
$\min_{x}w(x)=\alpha\cdot\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right)-(\alpha-1)D_{\alpha}(P||Q^{B}).$ (57)
We substitute this back into our violation condition (55) and obtain
$D_{\alpha}(P||Q^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right)-(\alpha-1)D_{\alpha}(P||Q^{B}),$ (58)
or equivalently,
$\alpha D_{\alpha}(P||Q^{B})<\alpha\cdot\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right).$ (59)
Again noticing that $\alpha<0$, the violation condition (55) reduces to
$D_{\alpha}(P||Q^{B})>\max_{x}\ln\left(\frac{P(x)}{Q^{B}(x)}\right),$ (60)
which is Eq. (15). Since this never holds, rationality is never violated.
Appendix G Derivation of the work bound
In this Appendix, we show the derivation of Eq. (23). Recall that the optimal distribution $Q^{A*,\mu}(x)$ is given by Eq. (22), for which we define $Z(\mu)=\sum_{x^{\prime}}P(x^{\prime})^{\mu}Q^{B}(x^{\prime})^{1-\mu}$. We can show that
$D(Q^{A*,\mu}\|Q^{B})=\mu\sum_{x}Q^{A*,\mu}\ln\frac{P(x)}{Q^{B}(x)}-\ln Z(\mu),$ (61)
$D(Q^{A*,\mu}\|P)=-(1-\mu)\sum_{x}Q^{A*,\mu}\ln\frac{P(x)}{Q^{B}(x)}-\ln Z(\mu).$ (62)
By noting that $D_{\mu}(P\|Q^{B})=-\frac{1}{1-\mu}\ln Z(\mu)$, we obtain
$D(Q^{A*,\mu}\|Q^{B})=D_{\mu}(P\|Q^{B})-\frac{\mu}{1-\mu}D(Q^{A*,\mu}\|P).$ (63)
Now recall that $Q^{A*,\mu}=\lambda_{x^{n}}$. Because of the risk constraint (21),
$D(Q^{A*,\mu}\|Q^{B})\geq D_{\mu}(P\|Q^{B})-\frac{\mu}{1-\mu}\frac{1}{n}\ln\frac{1}{\epsilon}.$ (64)
Going back to the work expression (18), we obtain
$W_{n}(x^{n})\geq k_{B}TnD(Q^{A*,\mu}||Q^{B})\geq k_{B}Tn\left(D_{\mu}(P\|Q^{B})-\frac{\mu}{1-\mu}\frac{1}{n}\ln\frac{1}{\epsilon}\right),$ (65)
which proves Eq. (23).